# Benchmarking and Improving Large Vision-Language Models for Fundamental Visual Graph Understanding and Reasoning

### v1

@@ 4 The MCDGraph Framework
To enhance the ability of open-source LVLMs to understand and reason on visual graphs, we propose MCDGraph, a self-supervised fine-tuning framework designed to enhance LVLMs’ ability to capture structural and relational information within visual graphs. As illustrated in Figure 5, MCDGraph comprises three key tasks: Masked Graph Infilling, Contrastive Graph Discrimination, and Graph Description.
@@
Table 5: Model performance (Acc/F1/EM_F1 for QA and F1 for FC) on various tasks. $\uparrow$ indicates an improvement compared to the original model. The complete experimental results are shown in Table 10 and 10.
@@ 4.1 Task1: Masked Graph Infilling
In Masked Graph Infilling, we randomly mask either nodes or edges in a visual graph and challenge the model to predict the masked element based on the partially observed graph and the corresponding text triples.
@@
where $M$ denotes the masked content in the graph, and $\hat{G}$, $I$, $T$ denotes the masked graph, task instruction, and text triples of the original graph, respectively. This task encourages LVLMs to infer missing structure information, improving its ability to understand graph structure and the relationships between elements.
Figure 5: Overall illustration of MCDGraph.
@@ 4.3 Task3: Graph Description
Graph Description task requires LVLMs to generate a textual description of a given visual graph. The description should include the total number of nodes and edges, as well as the names of all the nodes in the graph,
@@ 5.1 Experimental Setup
We validate the effectiveness of MCDGraph on top two performing LVLMs on VGCure, i.e., Qwen2-VL and InternVL2. We collect a new set of visual graphs beyond VGCure with synthetic structures and automatically create 20k training samples for MCDGraph. To prevent catastrophic forgetting, we apply LoRA (Hu et al., 2022) to efficiently enhance the LVLMs’ abilities while preserving their original performance. More details about training samples and implementation are available in Appendix C.
@@ 5.2 Results and Analysis
Main Results Table 5 compares the performance of Qwen2-VL and InternVL2 before and after applying MCDGraph on both visual graph understanding and reasoning tasks. It can be observed that MCDGraph improves the performance of LVLMs on almost all tasks, demonstrating the effectiveness of the proposed method. Notably, the improvement of LVLMs is particularly impressive on edge-related tasks, i.e., Edge Number Query (EN) and Nested Relation Query (NR), that are relatively difficult for LVLMs. This suggests that MCDGraph enhances LVLMs’ ability to understand the graph structures. In addition, although MDCGraph does not optimize for the tasks in VGCure, MCDGraph shows obvious improvement in both QA and FC samples for most tasks in VGCure. This demonstrates that the proposed method can improve the fundamental graph structure understanding of LVLMs, which leads to better performance on downstream tasks. We also present a case on NR to further understand the effectiveness of the proposed method, please refer to Appendix D for detailed information.
@@
Performance on Varying Complexity Figure 6 illustrates the performance improvements of MCDGraph on LVLMs across varying graph complexities on six representative tasks in VGCure. For both models, fine-tuning with MCDGraph consistently improves performance across most tasks and complexity levels. This indicates that MCDGraph effectively enhances the LVLMs’ ability to understand and reason over visual graphs. For simpler graphs, MCDGraph provides a smaller improvement where LVLMs already perform well. However, as graph complexity increases, the performance gap between fine-tuned and baseline models becomes more pronounced on most tasks, highlighting MCDGraph’s importance in handling more complex visual graphs. Due to the space limit, only the results with the number of edges are presented here, and the findings are similar in the other dimensions. More results are exhibited in Figures 24-29 in the Appendix.
@@ 6 Related Work
Multimodal Benchmark for Graphs
With the emergence of LVLMs, the evaluation of LVLMs for visual graph understanding and reasoning has gained increasing attention from researchers. Li et al. (2024d) and Wei et al. (2024) introduce VisionGraph and GVLQA based on existing NLGraph dataset (Li et al., 2024c), respectively, for evaluating the problem-solving capabilities of LVLMs in graph theory. Both of them contain a large number of synthetic visual graphs and the corresponding complex graph theory problems. Besides, Ai et al. (2024) propose a novel instruction-following benchmark for multimodal graph understanding and reasoning in both English and Chinese, which contains a number of real-world graph images with diverse structures across various domains. However, these benchmarks are proposed for challenging graph theory problems or multi-hop reasoning tasks where LVLMs achieve unsatisfactory results. The goal of our VGCure is to explore the fundamental understanding and reasoning abilities of LVLMs on visual graphs in order to explore the underlying reasons for the failure of LVLMs.
Boosting LVLMs for Visual Graph Reasoning
Aware of the limitations of current LVLMs on visual graph reasoning tasks, several works propose novel methods to improve the performance of LVLMs on visual graphs. Li et al. (2024d) present a Description-Program-Reasoning (DPR) chain to improve the logical accuracy in reasoning processes via graphical structure description generation and algorithm-aware multi-step reasoning. Wei et al. (2024) propose an end-to-end framework, GITA, to systematically integrate visual information into instruction-based graph reasoning. Furthermore, Deng et al. (2024) introduce GraphVis, which uses a curriculum fine-tuning scheme to train LVLMs on basic graphical feature recognition, followed by reasoning on visual graph QA tasks. Different from current methods which focus on optimizing LVLMs for specific downstream tasks, the proposed MCDGraph is a general-purpose, self-supervised method that can integrally improve the fundamental understanding and reasoning of LVLMs on visual graphs, thereby can be flexibly applied to most downstream graph-learning tasks.
@@ 7 Conclusion
In this paper, we first introduce VGCure, a comprehensive benchmark comprising 22 tasks designed to evaluate LVLMs’ fundamental understanding and reasoning capabilities on visual graphs. Experimental results on 14 LVLMs reveal prominent limitations of LVLMs on VGCure, particularly in capturing relational or structural information within graphs. Based on this observation, we further propose MCDGraph, a structure-aware self-supervised method designed to improve open-source LVLMs’ structural understanding of visual graphs. Experiments validate the effectiveness of MCDGraph, confirming its potential to improve LVLMs’ performance on fundamental visual graph learning tasks, as well as the robustness against complex graphs.
@@
  • Styles of Visual Graph. Although we introduce a variety of graph structures from both synthetic and real-world sources, we use the same visualization style, e.g., the same node shapes, colors, and layouts, for the generation of visual graphs. Thus, the single graph style might lead to biased evaluation results.
@@
  • Evaluation of MCDGraph. We only evaluate MCDGraph on the LVLMs’ fundamental visual graph understanding and reasoning abilities, lacking the validation in real-world scenarios, e.g., KGQA. In addition, due to resource limitations, we do not evaluate our method on more and larger LVLMs.
@@ A.2 Tasks Generation
To ensure the correctness of the generated samples, we first search for relevant paths in the given graph that satisfy the conditions of the task. Then the final QA samples and FC samples are generated based on the paths and corresponding templates. If no path exists, the generation of samples for the task is skipped. The final statistics of VGCure and the example samples with undirected graph for each task are shown in Table 6 and Table 7, respectively.
@@ B.2 Evaluation Metrics
For the QA samples of NQ, NN, CR, RA, SRN, and NR tasks, we use (macro-averaged) F1 score and Hits@1 as in the previous QA benchmarks Rajpurkar et al. (2016); Zhang et al. (2018). For the QA samples of CP task, we employ EM_F1, which iss the macro F1 score calculated based on the exact match between the predicted path and the ground truth path, and Label_Acc, which measures the accuracy of the model’s prediction on whether a path exists or not. For the QA samples of NNu, EN and DQ tasks, we compute the accuracy between predicted answers and ground truth. For the FC samples of all tasks, following Si et al. (2024a, b), we use macro F1 and accuracy as the metrics.
@@
Table 7: Examples with undirected graph for each task in VGCure. These samples all correspond to the SFN graph shown in Figure LABEL:fig:example_graphs(i).
Model Lora_rank Lora_ alpha Global Batch Size Learning rate Epoch
Qwen2-VL 8 16 64 1e-4 5
InternVL2 64 128 64 4e-5 1
Table 8: Hyperparameters for training
Appendix C Experimental Setup for Training
C.1 Training Samples
@@
C.2 Implementation Details
For Qwen2-VL, we employ the LoRA-based supervised fine-tuning scripts provided by LLaMA-Factory [^3] https://github.com/hiyouga/LLaMA-Factory. For InternVL2, we perform LoRA-based fine-tuning based on the code and documentation provided officially [^4] https://github.com/OpenGVLab/InternVL. The hyperparameters used for training are shown in Table 8. All the experiments are finished on 4 A100 GPUs with 80GB memory.
Appendix D Case Study
@@
Figure 17: Model performance (F1) on FC samples across various graph structures and tasks. Each radar chart corresponds to a specific question type, showing comparative results on QA samples across all graph types, where OF, PCQC and SN denotes Openflights, PubChemQC and Social network, respectively.

### v2

@@ 4 The MCDGraph Framework
To enhance the ability of LVLMs to understand and reason on visual graphs, we propose MCDGraph, a self-supervised fine-tuning framework designed to improve LVLMs’ ability to capture structural and relational information within visual graphs. As illustrated in Fig.5, MCDGraph comprises three key tasks: Masked Graph Infilling, Contrastive Graph Discrimination, and Graph Description.
Figure 5: Overall illustration of MCDGraph.
@@
Table 5: Model performance (Acc/F1/EM_F1 for QA and F1 for FC) on various tasks. $\uparrow$ indicates an improvement compared to the original model. The complete experimental results are shown in Tab.17 and 17.
@@ 4.1 Task1: Masked Graph Infilling
For this task, we randomly mask either nodes or edges in a visual graph and challenge the model to predict the masked element based on the partially observed graph. Since anonymized visual graphs contain no semantic information, we also provide the corresponding text triples as input.
@@
where $\hat{G}$, $I$, $T$ denotes the masked graph, task instruction, and text triples of the original graph, respectively. This task encourages LVLMs to infer missing structure information, improving their ability to understand graph structure and the relationships between elements.
@@ 4.3 Task3: Graph Description
Graph Description task requires LVLMs to generate a textual description of a given visual graph, including the total number of nodes and edges, as well as the names of all the nodes in the graph,
@@ 5.1 Experimental Setup
We validate the effectiveness of MCDGraph on top two performing LVLMs on VGCure, i.e., Qwen2-VL and InternVL2. We collect a new set of anonymized visual graphs beyond VGCure with synthetic structures and automatically create 20k training samples for MCDGraph. To prevent catastrophic forgetting, we apply LoRA (Hu et al., 2022) to efficiently enhance the LVLMs’ abilities while preserving their original performance. More details about training samples and implementation are available in Appendix D.
@@ 5.2 Results and Analysis
Main Results Tab.5 compares the performance of Qwen2-VL and InternVL2 before and after applying MCDGraph on both visual graph understanding and reasoning tasks. We can observe that (I) MCDGraph improves the performance of LVLMs on almost all tasks, demonstrating the effectiveness of the proposed method. (II) The improvement of LVLMs is particularly impressive on edge-related tasks, i.e., Edge Number Query (EN) and Nested Relation Query (NR), that are relatively difficult for LVLMs. This suggests that MCDGraph enhances LVLMs’ ability to capture structural information. (III) Although MCDGraph does not optimize for the tasks in VGCure, it still shows obvious improvement in both QA and FC samples for most tasks in VGCure. This demonstrates that the proposed method can improve the fundamental graph structure understanding capabilities of LVLMs, which leads to better performance on downstream tasks. We also present a case on NR to further understand the effectiveness of the proposed method, please refer to Appendix G for detailed information.
@@
Performance on Varying Complexity Fig.6 illustrates the performance improvements of MCDGraph on LVLMs across varying graph complexities on six representative tasks in VGCure. Fine-tuning with MCDGraph consistently improves performance across most tasks and complexity levels, demonstrating its its effectiveness in enhancing the LVLMs’ ability to understand and reason over visual graphs. For simpler graphs, the improvement is smaller, where LVLMs already perform well, but as complexity increases, the performance gap between fine-tuned and baseline models becomes more pronounced, highlighting MCDGraph’s importance in handling more complex visual graphs. Due to the space limit, the results for other dimensions are shown in Figs.31-36 in the Appendix, with conclusions similar to those above.
Generalization of MCDGraph To validate the generalization of our method, we regenerate 50 visual graphs with different visual styles and naming conventions of nodes and edges from those in VGCure for each graph structure. The details and results are presented in Appendix E.
Model VisionGraph FactKG
Connect Cycle MaxFlow Accuracy Precision Recall F1
Qwen2-VL 55.8 52.88 1.72 79.60 81.18 79.13 79.13
w MCDGraph 53.37 52.88 5.17 80.15 81.71 79.69 79.71
InternVL2 46.9 52.88 6.9 79.41 80.93 78.95 78.95
w MCDGraph 54.72 52.88 8.62 79.78 81.34 79.32 79.33
Table 6: Model performance on downstream tasks.
5.3 Results on Downstream Reasoning Tasks
To further demonstrate the scalability and applicability of our method, we evaluate MCDGraph on graph-related downstream reasoning tasks.
VisionGraph We first evaluate the performance of MCDGraph on three representative graph theory problems in VisionGraph Li et al. (2024c). As the results shown in Tab.6, MCDGraph generally improve the performance of LVLMs on these tasks, especially for the relatively difficult Maximum Flow task. This confirms the effectiveness and scalability of our method.
FactKG We then evaluate MCDGraph on FactKG Kim et al. (2023b), a knowledge graph-based fact verification dataset collected from real-world data. As Tab.6 shows, MCDGraph gives consistently better performance than LVLMs on FactKG, suggesting that the proposed method can improve LVLMs in real-world graph-related tasks. Note that FactKG not only requires fundamental visual graph understanding and reasoning abilities but also relies on LVLMs’ understanding of semantic and logical relationships between entities, which our method does not address. Therefore, the improvement of LVLMs is not as significant as that of the graph theory problems. We also evaluate our method on two normal VQA tasks and the results are available in Appendix F.
@@ 6 Related Work
Multimodal Benchmark for Graphs Li et al. (2024c) and Wei et al. (2024) introduce VisionGraph and GVLQA, respectively, for testing the problem-solving capabilities of LVLMs in graph theory. Both of them contain numerous synthetic visual graphs and complex graph theory problems. Besides, Ai et al. (2024) propose a novel instruction-following benchmark for multimodal graph understanding and reasoning, which contains a number of real-world graph images with diverse structures across various domains. However, these benchmarks focus on specific downstream tasks where LVLMs perform poorly. The goal of VGCure is to assess LVLMs’ fundamental understanding and reasoning abilities on visual graphs to identify the reasons for their failures.
Boosting LVLMs for Visual Graph Reasoning Li et al. (2024c) propose a Description-Program-Reasoning (DPR) chain to enhance logical accuracy through graphical structure description and multi-step reasoning. Wei et al. (2024) introduce GITA, an end-to-end framework integrating visual information into instruction-based graph reasoning. Additionally, Deng et al. (2024) present GraphVis, which uses curriculum fine-tuning for training LVLMs on feature recognition and visual graph QA tasks. Unlike current methods, our MCDGraph is a general-purpose, self-supervised approach that improves the fundamental understanding and reasoning of LVLMs on visual graphs, making it adaptable to most graph-related downstream tasks.
@@ 7 Conclusion
This paper introduces VGCure, a comprehensive benchmark comprising 22 tasks to evaluate LVLMs’ fundamental understanding and reasoning capabilities on visual graphs. Experiments on 14 LVLMs reveal significant limitations, especially in capturing structural information. To this end, we propose MCDGraph, a structure-aware self-supervised method to enhance open-source LVLMs’ structure learning abilities. Extensive experiments validate the effectiveness of our method across a wide range of graph-related tasks.
@@
  • Experiments on Larger LVLMs. Due to limited resources, the majority of our experiments are performed only on LVLMs with around 7B parameters, lacking performance evaluation and improvement of larger models with more parameters.
@@ A.1 Graphs Generation
For all the above structures except SBM, the number of nodes during generation is an arbitrary integer in the range $[7,15]$.
@@ A.2 Tasks Generation
To ensure the correctness of the generated samples, we first search for relevant paths in the given graph that satisfy the conditions of the task. Then the final QA samples and FC samples are generated based on the paths and corresponding templates. If no path exists, the generation of samples for the task is skipped. The final statistics of VGCure and the example samples with undirected graph for each task are shown in Tab.7 and Tab.8, respectively.
@@ B.2 Evaluation Metrics
For the QA samples of NQ, NN, CR, RA, SRN, and NR tasks, we use (macro-averaged) F1 score and Hits@1 as in the previous QA benchmarks Rajpurkar et al. (2016); Zhang et al. (2018). For the QA samples of CP task, we employ EM_F1, which is the macro F1 score calculated based on the exact match between the predicted path and the ground truth path, and Label_Acc, which measures the accuracy of the model’s prediction on whether a path exists or not. For the QA samples of NNu, EN and DQ tasks, we compute the accuracy between predicted answers and ground truth. For the FC samples of all tasks, following Zhu et al. (2023) and Si et al. (2024) , we use macro F1 and accuracy as the metrics.
@@
Table 8: Examples with undirected graph for each task in VGCure. These samples all correspond to the SFN graph shown in Fig.LABEL:fig:example_graphs(i).
Error Type QA samples Visual Graph
Relation Misunderstanding
Q: What is the relation from the entity that is R5 of E3 to E1?
A: [R1]
P: [R6]
Fig.LABEL:fig:error_graphs(a)
Complexity Limitation
Q: What is the relation from the entity that is R7 of E1 to E7?
A: [R5]
P: [R9]
Fig.LABEL:fig:error_graphs(b)
Structural Hallucination
Q: Is there a path between E1 and E3?
A: Yes. The shortest paths are [[E1, E6, E3]]
P: Yes. The shortest paths are [E1, E3] and [E1, E4, E3].
Fig.LABEL:fig:error_graphs(c)
Question-based Answering
Q: Which entities are R5 of E2 as well as R1 of E1?
A: [E11]
P: [E2, E1]
Fig.LABEL:fig:error_graphs(b)
Off-target Generation
Q: What is the relation between E2 and the entity that is connected to E4 via R9?
A: [R2]
P: [Edge, Node]
Fig.LABEL:fig:error_graphs(d)
Format Error
Q: Is there a path between E4 and E1?
A: Yes. The shortest paths are [[E4, E3, E2, E1]]
P: Yes. The shortest paths are [[E4, R9, E3, R2, E2, R6, E1]].
Fig.LABEL:fig:error_graphs(d)
Table 9: Examples of each error type, where Q denotes question, A denotes gold answer and P denotes prediction.
Appendix C Why LVLMs Fail on Fundamental Tasks?
C.1 Comparison with Backbone LLMs
To compare the performance of the LVLMs with the corresponding LLMs on the VGCure task, we first randomly select 1000 samples for each task and convert the corresponding visual graphs into text triples to construct the text version of the samples, and then evaluate them with the corresponding backbone LLMs.
The complete prompts for QA samples are as follows:
[…]
The complete prompt for FC samples are as follow:
[…]
The Connective Path Query (CP) task is ignored here due to LLMs’ poor instruction-following ability on this task.
C.2 Error Analysis
We randomly select a batch of results generated by Qwen2-VL and InternVL2 on QA samples in VGCure and attribute their errors to the following six types:
  • Relation Misunderstanding: Failure to properly understand the relations between entities, leading to erroneous reasoning.
  • Complexity Limitation: When faced with intricate or highly complex graph structures, LVLMs struggle to process and understand the information effectively, often resulting in incomplete or erroneous outputs.
  • Structural Hallucination: Generating or perceiving structures that do not exist, leading to erroneous or misleading results that do not match the actual visual graph.
  • Question-based Answering: Directly using the entities mentioned in the question as answers, thus bypassing the need for deeper relation understanding or logical reasoning in the visual graph.
  • Off-target Generation: Deviation from the original task or misunderstanding of the question, leading to the generation of irrelevant answers.
  • Format Error: The output generated by the model is incorrectly formatted or unexpected.
Examples of each error are shown in the Tab.9.
Appendix D Experimental Setup for Training
D.1 Training Samples
@@
Model Lora_rank Lora_ alpha Global Batch Size Learning rate Epoch
Qwen2-VL 8 16 64 1e-4 5
InternVL2 64 128 64 4e-5 1
Table 10: Hyperparameters for training
D.2 Implementation Details
For Qwen2-VL, we employ the LoRA-based supervised fine-tuning scripts provided by LLaMA-Factory [^3] https://github.com/hiyouga/LLaMA-Factory. For InternVL2, we perform LoRA-based fine-tuning based on the code and documentation provided officially [^4] https://github.com/OpenGVLab/InternVL. The hyperparameters used for training are shown in Tab.10. All the experiments are finished on 4 A100 GPUs with 80GB memory.
Models ScienceQA AOKVQA
Qwen2-VL 85.13 84.37
w MCDGraph 81.31 83.67
InternVL2 97.07 85.41
w MCDGraph 96.88 84.45
Table 11: Model performance (Acc) on normal VQA tasks.
Appendix E Generalization of MCDGraph
E.1 Impact on Visual Styles
To validate the generalization of our method, we employ NetworkX and Matplotlib to regenerate 50 visual graphs with different visual styles from those in VGCure for each graph structure to explore the impact of visual graph styles. The examples of the same visual graph with different styles are illustrated in Fig.21 and Fig.22.
As the results shown in Tabs.12 and 13, although LVLMs never encounter the different style of visual graph during fine-tuning, our method can still improve the performance of the LVLMs on almost all tasks. This demonstrates the ability of our method to enhance LVLMs’ ability in capturing the structural information in visual graphs with excellent generalization. In addition, compared to Tab.5, it can be noticed that the experimental results before and after the change of style are similar, which indicates that the style itself has no effect on the evaluation of LVLMs.
E.2 Impact of Naming Conventions
To explore the impact of altering the naming conventions, we regenerated the visual graphs and samples using the following new naming conventions.
[…]
  • Name-R: The nodes in the visual graph are renamed to a simple and common human name without any semantic bias, like “John”, “Jane”, “Mike”, “Mary”, etc. The names of the edges remain as they are, i.e., R1, R2, …, R10.
The results are shown in Tabs.14 and 15. We can observe that after altering the naming conventions, LVLMs continue to show similar trends on most of tasks. Therefore, the experimental analysis in the main text still holds. Meanwhile, the performance of LVLMs decreases on most of the tasks when confronted with different names. This might be due to the fact that the new naming conventions gives longer names to nodes and edges and recognizing these information increases the difficulty of the task. It is worth noting that in the face of the new naming conventions, our MCDGraph still improves the performance of LVLMs on most tasks, although the fine-tuning still uses the original naming conventions. This strongly demonstrates the effectiveness as well as the generalization of our method.
Appendix F Results on Normal VQA Tasks
We evaluate our MCDGraph on two normal VQA tasks, i.e., ScienceQA Lu et al. (2022) and AOKVQA Schwenk et al. (2022). According to the results shown in Tab.11, we can observe that after MCDGraph, LVLMs perform worse on these two VQA tasks. This is due to the fact that our MCDGraph is proposed for visual graph understanding and reasoning tasks, which are very different from normal VQA tasks. And it is worth noting that our method does not overly compromise the LVLMs’ performance on normal VQA tasks, which we consider acceptable
@@ Appendix F Results on Normal VQA Tasks
Appendix G Case Study
Fig.20 illustrates a case on Nested Relation Query task. We can observe that Qwen2-VL, LLaVA-13B and mPLUG-Owl3 all make the error of Question-based Answering when confronted with this question, i.e., using “R1”, “E9”, and “E3” mentioned in the question as the generated answer. InternLM-XC2.5 makes the error of Relation Misunderstanding, i.e., the edge pointing from E3 to E1 is also used as the answer. However, Qwen2-VL after applying MCDGraph can answer this question correctly. This demonstrates that the proposed method can improve the fundamental graph structure understanding of LVLMs, thus avoiding the occurrence of the previously mentioned errors.
Figure 21: An example of visual graph with different style for the experimental results in Tab.12
Figure 22: An example of visual graph with different style for the experimental results in Tab.13
Models Understanding Reasoning
NNu EN DC DQ NQ NN CR CN RA SRN NR CP
QA Samples
Qwen2-VL 98.20 18.10 - 50.70 18.64 15.66 22.66 - 14.83 20.45 16.95 11.04
w MCDGraph 99.40 $\uparrow$ 40.20 $\uparrow$ - 56.80 $\uparrow$ 28.44 $\uparrow$ 14.64 27.02 - 13.69 $2$uparrow 14.73 11.99 $\uparrow$
InternVL2 73.70 8.60 - 48.40 26.33 18.26 26.19 - 9.85 21.75 11.20 1.60
w MCDGraph 95.80 $\uparrow$ 35.70 $\uparrow$ - 52.90 $\uparrow$ 28.89 $\uparrow$ 19.90 $\uparrow$ 29.24 $\uparrow$ - 11.73 $\uparrow$ 23.23 $\uparrow$ 18.04 $\uparrow$ $\uparrow$ 14.49
FC Samples
Qwen2-VL 71.37 61.78 86.53 73.82 73.03 48.00 77.35 39.98 36.04 60.22 45.31 42.93
w MCDGraph 88.45 $\uparrow$ 63.14 $\uparrow$ 94.43 $\uparrow$ 76.26 $\uparrow$ 78.59 $\uparrow$ 63.17 $\uparrow$ 82.63 $\uparrow$ 53.78 $\uparrow$ 54.33 $\uparrow$ 66.26 $\uparrow$ 65.95 $\uparrow$ 56.30 $\uparrow$
InternVL2 75.43 34.80 86.85 63.68 74.52 61.12 74.23 47.22 33.48 55.85 55.32 53.15
w MCDGraph 83.82 $\uparrow$ 70.14 $\uparrow$ 83.48 60.77 79.69 $\uparrow$ 71.39 $\uparrow$ 78.30 $\uparrow$ 52.38 $\uparrow$ 51.24 $\uparrow$ 59.83 $\uparrow$ 68.58 $\uparrow$ 44.73
Table 12: Model performance (Acc/F1/EM_F1 for QA and F1 for FC) on various tasks with different visual styles. The example of the corresponding visual style is shown in Fig.21
Models Understanding Reasoning
NNu EN DC DQ NQ NN CR CN RA SRN NR CP
QA Samples
Qwen2-VL 95.30 20.80 - 40.00 13.95 8.73 19.04 - 11.73 14.89 13.16 3.17
w MCDGraph 98.80 $\uparrow$ 17.30 - 50.50 $\uparrow$ 22.33 $\uparrow$ 11.56 $\uparrow$ 21.92 $\uparrow$ - 11.05 17.84 $\uparrow$ 11.60 0.12
InternVL2 85.20 17.90 - 39.60 24.21 19.99 22.26 - 8.06 19.46 12.03 2.25
w MCDGraph 97.20 $\uparrow$ 35.70 $\uparrow$ - 43.00 $\uparrow$ 26.46 $\uparrow$ 19.12 23.54 $\uparrow$ - 11.00 $\uparrow$ 20.45 $\uparrow$ 14.80 $\uparrow$ 12.11 $\uparrow$
FC Samples
Qwen2-VL 70.88 68.06 66.94 57.90 65.09 43.62 67.32 34.28 34.79 53.23 42.54 42.89
w MCDGraph 76.36 $\uparrow$ 58.47 85.20 $\uparrow$ 74.38 $\uparrow$ 73.69 $\uparrow$ 56.97 $\uparrow$ 75.05 $\uparrow$ 48.68 $\uparrow$ 53.84 $\uparrow$ 60.99 $\uparrow$ 60.44 $\uparrow$ 56.93 $\uparrow$
InternVL2 68.42 38.35 91.84 64.58 65.45 62.91 72.42 48.17 34.60 57.88 57.77 52.48
w MCDGraph 69.45 $\uparrow$ 69.11 $\uparrow$ 84.10 62.09 74.68 $\uparrow$ 71.52 $\uparrow$ 75.19 $\uparrow$ 52.97 $\uparrow$ 50.89 $\uparrow$ 61.00 $\uparrow$ 66.44 $\uparrow$ 43.19
Table 13: Model performance (Acc/F1/EM_F1 for QA and F1 for FC) on various tasks with different visual styles. The example of the corresponding visual style is shown in Fig.22
Models Understanding Reasoning
NNu EN DC DQ NQ NN CR CN RA SRN NR CP
QA Samples
Qwen2-VL 94.30 18.10 - 48.40 17.55 11.34 13.72 - 10.10 11.82 5.34 10.63
w MCDGraph 94.10 16.00 - 63.50 $\uparrow$ 17.25 8.50 14.56 $\uparrow$ - 8.29 12.81 $\uparrow$ 12.71 $\uparrow$ 13.00 $\uparrow$
InternVL2 65.61 19.15 - 47.94 17.45 12.11 15.75 - 8.87 13.54 8.25 8.12
w MCDGraph 89.30 $\uparrow$ 34.60 $\uparrow$ - 49.20 $\uparrow$ 17.59 $\uparrow$ 12.88 $\uparrow$ 14.63 - 8.29 13.75 $\uparrow$ 4.52 5.45
FC Samples
Qwen2-VL 74.18 66.99 91.90 58.77 58.34 37.99 59.80 34.25 34.51 45.99 35.39 39.40
w MCDGraph 85.24 $\uparrow$ 66.42 90.93 81.77 $\uparrow$ 77.30 $\uparrow$ 56.79 $\uparrow$ 76.47 $\uparrow$ 53.06 $\uparrow$ 52.85 $\uparrow$ 59.45 $\uparrow$ 58.51 $\uparrow$ 58.72 $\uparrow$
InternVL2 72.82 41.92 96.19 65.34 73.36 61.51 77.15 52.21 34.79 55.23 57.50 53.17
w MCDGraph 73.86 $\uparrow$ 72.70 $\uparrow$ 94.28 59.99 78.23 $\uparrow$ 69.02 $\uparrow$ 74.32 52.12 54.71 $\uparrow$ 57.41 $\uparrow$ 68.51 $\uparrow$ 44.20
Table 14: Model performance (Acc/F1/EM_F1 for QA and F1 for FC) on various tasks with Node-Edge naming convention.
Models Understanding Reasoning
NNu EN DC DQ NQ NN CR CN RA SRN NR CP
QA Samples
Qwen2-VL 21.00 22.90 - 51.10 17.78 11.80 18.31 - 11.50 15.33 8.66 14.02
w MCDGraph 13.50 25.70 $\uparrow$ - 63.30 $\uparrow$ 18.66 $\uparrow$ 9.56 18.91 $\uparrow$ - 10.13 15.90 $\uparrow$ 10.81 $\uparrow$ 11.72
InternVL2 32.80 17.70 - 55.10 19.77 14.61 21.60 - 10.69 16.20 11.29 11.79
w MCDGraph 24.40 36.00 $\uparrow$ - 54.20 21.10 $\uparrow$ 13.32 21.44 - 8.15 16.01 17.77 $\uparrow$ 11.84 $\uparrow$
FC Samples
Qwen2-VL 86.35 67.03 93.73 69.39 74.05 41.30 71.54 43.59 38.54 58.05 42.43 44.18
w MCDGraph 68.68 69.10 $\uparrow$ 96.45 $\uparrow$ 78.20 $\uparrow$ 76.96 $\uparrow$ 57.61 $\uparrow$ 82.66 $\uparrow$ 56.07 $\uparrow$ 55.19 $\uparrow$ 63.51 $\uparrow$ 63.17 $\uparrow$ 56.45 $\uparrow$
InternVL2 69.34 39.80 87.75 62.98 76.17 69.06 72.14 57.15 39.82 56.87 65.83 48.31
w MCDGraph 78.05 $\uparrow$ 71.65 $\uparrow$ 75.85 55.84 81.23 $\uparrow$ 71.50 $\uparrow$ 79.29 $\uparrow$ 52.43 57.76 $\uparrow$ 53.85 67.60 $\uparrow$ 41.99
Table 15: Model performance (Acc/F1/EM_F1 for QA and F1 for FC) on various tasks with Name-R naming convention.
Figure 23: Model performance (Acc/F1) on QA samples across various graph structures and tasks.
Figure 24: Model performance (F1) on FC samples across various graph structures and tasks.